# Dynamics of the general $Q$-tensor model interacting with a rigid body

### v1

@@ 8. Global strong well-posedness for small data
First, we verify Theorem 5.2. We have already argued that $\Psi\colon\tilde{\mathcal{K}}_{\varepsilon}\to\tilde{\mathcal{K}}$ is well-defined. For $\varepsilon\in(0,\varepsilon_{0})$, with $\varepsilon_{0}>0$ as above, let $(z,p)$, $(z_{1},p_{1})$, $(z_{2},p_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$. By $\|(v_{0},Q_{0},\ell_{0},\omega_{0})\|_{\mathrm{X}_{\gamma}}\leq\delta$, and exploiting Theorem 6.1(b) as well as Section 8, Section 8 and Section 8, we find
$\|\Psi(z,p)\|_{\tilde{\mathcal{K}}}\leq C_{\mathrm{MR}}(\delta+C\varepsilon^{2})\enspace\text{and}\enspace\|\Psi(z_{1},p_{1})-\Psi(z_{2},p_{2})\|_{\tilde{\mathcal{K}}}\leq C_{\mathrm{MR}}C\varepsilon.$
With $\delta\leq\nicefrac{{1}}{{2C_{\mathrm{MR}}}}$ and $\varepsilon\leq\min\{\varepsilon_{0},\nicefrac{{1}}{{2C_{\mathrm{MR}}C}},\nicefrac{{1}}{{2C}},\sqrt{\nicefrac{{\delta}}{{C}}}\}$, we conclude the self-map and contraction property of $\Psi$. Thus, there exists a unique fixed point $(z_{*},p_{*})=(v_{*},Q_{*},\ell_{*},\omega_{*},p_{*})$, being the unique global strong solution to (5.11)–(5.13). Theorem 5.2 then is a consequence of the definition of $\tilde{\mathcal{K}}$ in (8.3).
The last task is to deduce Theorem 4.2 from Theorem 5.2. The latter implies $(\ell,\omega)\in\mathrm{W}^{1,p}(0,\infty)^{6}$, so $h^{\prime}$, $\Omega$ and $X$ can be obtained as indicated in Section 5. For $v$, $Q$ and $p$, we invoke the backward change of coordinates from Section 5. By construction, this is the solution to (2.5)–(2.9) in the asserted regularity class. Finally, we prove that $\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}$ for all $t\in[0,\infty)$. The shape of $\mathcal{S}(t)$ at time $t$ as introduced in Section 2 and the above estimates imply that for all $t\in[0,\infty)$, we have
$\mathrm{dist}(\mathcal{S}(t),\mathcal{S}_{0})\leq|h(t)|+|\mathbb{O}(t)-\operatorname{Id}_{3}|<\nicefrac{{r}}{{2}},\enspace\text{so}\enspace\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}\enspace\text{by}\enspace\mathrm{dist}(\mathcal{S}_{0},\mathcal{O})>r.\qed$
Acknowledgements This research is supported by the Basque Government through the BERC 2022-2025 program and by the Spanish State Research Agency through BCAM Severo Ochoa excellence accreditation CEX2021-01142-S funded by MICIU/AEI/10.13039/501100011033 and through Grant PID2023-146764NB-I00 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union. Felix Brandt and Matthias Hieber acknowledge the support by DFG project FOR 5528. Arnab Roy would like to thank the Alexander von Humboldt-Foundation, Grant RYC2022-036183-I funded by MICIU/AEI/10.13039/501100011033 and by ESF+.

### v2

@@ 8. Global strong well-posedness for small data
First, we verify Theorem 5.2. We have already argued that $\Psi\colon\tilde{\mathcal{K}}_{\varepsilon}\to\tilde{\mathcal{K}}$ is well-defined. For $\varepsilon\in(0,\varepsilon_{0})$, with $\varepsilon_{0}>0$ as above, let $(z,\pi)$, $(z_{1},\pi_{1})$, $(z_{2},\pi_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$. By $\|(v_{0},Q_{0},\ell_{0},\omega_{0})\|_{\mathrm{X}_{\gamma}}\leq\delta$, and exploiting Theorem 6.1(b) as well as Section 8, Section 8 and Section 8, we find
$\|\Psi(z,\pi)\|_{\tilde{\mathcal{K}}}\leq C_{\mathrm{MR}}(\delta+C\varepsilon^{2})\kern 5.0pt\text{and}\kern 5.0pt\|\Psi(z_{1},\pi_{1})-\Psi(z_{2},\pi_{2})\|_{\tilde{\mathcal{K}}}\leq C_{\mathrm{MR}}C\varepsilon.$
With $\delta\leq\nicefrac{{1}}{{2C_{\mathrm{MR}}}}$ and $\varepsilon\leq\min\{\varepsilon_{0},\nicefrac{{1}}{{2C_{\mathrm{MR}}C}},\nicefrac{{1}}{{2C}},\sqrt{\nicefrac{{\delta}}{{C}}}\}$, we conclude the self-map and contraction property of $\Psi$. Thus, there exists a unique fixed point $(z_{*},\pi_{*})=(v_{*},Q_{*},\ell_{*},\omega_{*},\pi_{*})$, being the unique global strong solution to (5.11)–(5.13). Theorem 5.2 then is a consequence of the definition of $\tilde{\mathcal{K}}$ in (8.3).
The last task is to deduce Theorem 4.2 from Theorem 5.2. The latter implies $(\ell,\omega)\in\mathrm{W}^{1,p}(0,\infty)^{6}$, so $h^{\prime}$, $\Omega$ and $X$ can be obtained as indicated in Section 5. For $v$, $Q$ and $\pi$, we invoke the backward change of coordinates from Section 5. By construction, this is the solution to (2.5)–(2.9) in the asserted regularity class. Finally, we prove that $\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}$ for all $t\in[0,\infty)$. The shape of $\mathcal{S}(t)$ at time $t$ as introduced in Section 2 and the above estimates imply that for all $t\in[0,\infty)$, we have
$\mathrm{dist}(\mathcal{S}(t),\mathcal{S}_{0})\leq|h(t)|+|\mathbb{O}(t)-\operatorname{Id}_{3}|<\nicefrac{{r}}{{2}},\kern 5.0pt\text{so}\kern 5.0pt\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}\kern 5.0pt\text{by}\kern 5.0pt\mathrm{dist}(\mathcal{S}_{0},\mathcal{O})>r.\qed$
Acknowledgements This research is supported by the Basque Government through the BERC 2022-2025 program and by the Spanish State Research Agency through BCAM Severo Ochoa excellence accreditation CEX2021-01142-S funded by MICIU/AEI/10.13039/501100011033 and through Grant PID2023-146764NB-I00 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union. Felix Brandt would like to express his gratitude to the German National Academy of Sciences Leopoldina for support through the Leopoldina Fellowship Program with grant number LPDS 2024-07. Felix Brandt and Matthias Hieber acknowledge the support by DFG project FOR 5528. Arnab Roy would like to thank the Alexander von Humboldt-Foundation, Grant RYC2022-036183-I funded by MICIU/AEI/10.13039/501100011033 and by ESF+. The authors would also like to thank the anonymous referees for the careful reading and valuable comments that helped improve the quality of the manuscript.